# Supplement

### v3

@@ 5.3 Bayesian MFPCA of individual CGM curves
Figure 7 provides the estimated variability (left panels) and percent variability explained (right panels) within level and diet. The first row of panels corresponds to the subject (level 1) while the second row of panels corresponds to the meal (level 2). Each color corresponds to a diet. As in the single level analyses, Figure 6 indicates qualitative similarity of estimates $\widehat{\phi}_{k}^{(1)}(t),\widehat{\phi}_{l}^{(2)}(t)$ across diets. The primary mode of variability within each diet is $\widehat{\phi}^{(1)}_{1}(t)$, a nearly constant function. Higher subject scores on this component correspond to higher subject-specific glucose over the entire observation window. The sample-to-sample variability in the shape of $\phi^{(1)}_{1}(t)$ is consistently low, indicating confidence in this functional form. There is some heterogeneity in the percent variability explained by this component, but it is $\geq 35$% for every diet. The first eigenfunction at the second level, $\widehat{\phi}^{(2)}_{1}(t)$, is also nearly constant with a small curvature and low sampling variability. This FPC explains between $25-33$% of the total variability according to Figure 7, indicating that a constant shift is the main source of variation in glucose responses both among and within participants. The second eigenfunction at the meal level, $\widehat{\phi}^{(2)}_{2}(t)$, explains $10-25$% of the total variability and has an $S$ pattern. As in the single-level case, larger scores on $\widehat{\phi}^{(2)}_{2}(t)$ correspond to lower pre- and immediately post-prandial glucose, then increasingly larger response one to four hours after the meal. The eigenfunctions $\widehat{\phi}^{(1)}_{2}(t),\widehat{\phi}^{(2)}_{3}(t)$ capture higher-order glycemic reaction patterns and explain much smaller proportions of variability ($\leq 10$%).
@@
Each FAST MFPCA model takes $\approx 6$ minutes to run $5000$ iterations on a standard laptop, with the first $2500$ discarded as burn-in. We again utilize the routine described in Section 2.5 to assess convergence of the scores and FPCs at both levels. This requires aligning each level separately. Final Gelman-Rubin statistics were $\leq 1.02$.
@@ 6 Discussion
Standard implementations of FPCA use a two-step procedure: (1) diagonalize a smooth estimate of the observed data covariance; and (2) conduct inference conditional on the FPC estimates from the first step. This does not account for sampling variability in the FPCs, which could be quite large for low to moderate signal-to-noise. Our approach is a single step joint Bayesian model that treats PCs as parameters in the model, which combines the orthogonal spline expansion from Ye, (2024), the polar decomposition-based parameter expansion from Jauch et al., (2021), and stabilizing constraints to efficiently jointly model all FPCA model components. Combining the right ingredients led to FAST, a stable, replicable, and easy to implement methodology. The central element of FAST is expressing the FPCs in terms of an orthonormal spline basis. This is different from the approach in Ye, (2024), as this spline basis: (1) is orthonormal in $L^{2}([0,1])$, rather than in a vector sense; and (2) retains the computational efficiency of B-splines (Liu et al.,, 2020). This spline expansion simplifies the problem of sampling the high dimensional FPCs to sampling a low dimensional spline coefficients. The FPC spline coefficient matrix must be orthonormal to produce the corresponding FPC samples. This is done using the polar decomposition parameter expansion of Jauch et al., (2021), applying the technique to the coefficient matrix rather than the data level FPCs.
@@
Many questions remain, including: (1) generalizability to non-Gaussian, multivariate, sparse, and longitudinal functional data; (2) theoretical guarantees and non-uniform priors on the FPC parameter Stiefel manifold; (3) eigenbasis dimension estimation; and (4) further computational improvements, possibly through adaptive sampling.
@@
[…]
$=\theta^{t}(\alpha\mathbf{P}_{0}+(1-\alpha)\mathbf{P}_{2})\theta$

### v4

@@ 5.3 Bayesian MFPCA of individual CGM curves
Figure 7 provides the estimated variability (left panels) and percent variability explained (right panels) within level and diet. The first row of panels corresponds to the subject (level 1) while the second row of panels corresponds to the meal (level 2). Each color corresponds to a diet. As in the single level analyses, Figure 6 indicates qualitative similarity of estimates $\widehat{\phi}_{k}^{(1)}(t),\widehat{\phi}_{l}^{(2)}(t)$ across diets. The primary mode of variability within each diet is $\widehat{\phi}^{(1)}_{1}(t)$, a nearly constant function. Higher subject scores on this component correspond to higher subject-specific glucose over the entire observation window. The sample-to-sample variability in the shape of $\phi^{(1)}_{1}(t)$ is consistently low, indicating confidence in this functional form. There is some heterogeneity in the percent variability explained by this component, but it is $\geq 30$% for every diet. The first eigenfunction at the second level, $\widehat{\phi}^{(2)}_{1}(t)$, is also nearly constant with a small curvature and low sampling variability. This FPC explains between 253525-35% of the total variability according to Figure 7, indicating that a constant shift is the main source of variation in glucose responses both among and within participants. The second eigenfunction at the meal level, $\widehat{\phi}^{(2)}_{2}(t)$, explains 102510-25% of the total variability and has an $S$ pattern. As in the single-level case, larger scores on $\widehat{\phi}^{(2)}_{2}(t)$ correspond to lower pre- and immediately post-prandial glucose, then increasingly larger response one to four hours after the meal. The eigenfunctions $\widehat{\phi}^{(1)}_{2}(t),\widehat{\phi}^{(2)}_{3}(t)$ capture higher-order glycemic reaction patterns and explain much smaller proportions of variability ($\leq 10$%).
@@
Each FAST MFPCA model takes 131-3 minutes to run $2000$ iterations on a standard laptop, with the first $1000$ discarded as burn-in. We again utilize the routine described in Section 2.5 to assess convergence of the scores and FPCs at both levels. This requires aligning each level separately. Final Gelman-Rubin statistics were <1.05<1.05.
@@ 6 Discussion
Standard implementations of FPCA use a two-step procedure: (1) diagonalize a smooth estimate of the observed data covariance; and (2) conduct inference conditional on the FPC estimates from the first step. This does not account for sampling variability in the FPCs, which could be quite large for low to moderate signal-to-noise. Our approach is a single-step joint Bayesian model that treats PCs as parameters in the model, which combines the orthogonal spline expansion from Ye, (2024), the parameter expansion based on polar decomposition from Jauch et al., (2021), and stabilizing constraints to efficiently jointly model all FPCA model components. Combining the right ingredients led to FAST, a stable, replicable, and easy to implement methodology. The central element of FAST is expressing the FPCs in terms of an orthonormal spline basis. This is different from the approach in Ye, (2024), as this spline basis: (1) is orthonormal in $L^{2}([0,1])$, rather than in a vector sense; and (2) retains the computational efficiency of B-splines (Liu et al.,, 2020). This spline expansion simplifies the problem of sampling the high dimensional FPCs to sampling a low dimensional spline coefficients. The FPC spline coefficient matrix must be orthonormal to produce the corresponding FPC samples. This is accomplished using the polar decomposition parameter expansion of Jauch et al., (2021), applying the technique to the coefficient matrix rather than the data level FPCs.
@@
Recall we chose a posterior addition which mixes second- and zero-order penalties of the form $\alpha\int f^{2}(t)dt+(1-\alpha)\int\{f^{\prime\prime}(t)\}^{2}dt$. For spline parameters $\theta$ such that $f(t)\approx\mathbf{B}(t)\theta$, there exist unique penalty matrices $\mathbf{P}_{0},\mathbf{P}_{2}$ such that $\int f^{2}(t)dt\approx\theta^{t}\mathbf{P}_{0}\theta$ and $\int\{f^{\prime\prime}(t)\}^{2}dt\approx\theta^{t}\mathbf{P}_{2}\theta$ (Craven and Wahba,, 1979; Kimeldorf and Wahba,, 1970; O’Sullivan,, 1986; Wahba,, 1983).
We proceed by defining both of the penalty components $\mathbf{P}_{0},\mathbf{P}_{2}$ separately. Let $\mathbf{B}(t)=[b_{1}(t)|\ldots,|b_{Q}(t)]$ represent the basis functions, chosen to be orthonormal in $L^{2}([0,1])$. First, we define the zero-order penalty $\mathbf{P}_{0}$ element-wise. This derivation leverages the orthonormal definition of the $b_{i}(t)$.
$(\mathbf{P}_{0})_{ij}$ $=\int_{0}^{1}\int_{0}^{1}b_{i}(t)b_{j}(t)dt$
$=\begin{cases}1\text{ when }i=j\\ 0\text{ otherwise}\end{cases}$
The resulting matrix $\mathbf{P}_{0}=\mathbf{I}_{Q}$ by the definition of the basis $\mathbf{B}(t)$.
$(\mathbf{P}_{2})_{ij}$ $=\int_{0}^{1}\int_{0}^{1}b^{\prime\prime}_{i}(t)b_{j}^{\prime\prime}(t)dt$
We approximate these quantities using numerical integration.
To ensure uniformity of penalization scale between $\mathbf{P}_{0}$ and $\mathbf{P}_{2}$, we scale both raw penalties such that their leading eigenvalues are 1. We then follow Goldsmith et al., (2015) and define $\mathbf{P}_{\alpha}=\alpha\mathbf{P}_{0}+(1-\alpha)\mathbf{P}_{2}$. This allows us to write the final penalty using a single quadratic form:
$\alpha\theta^{t}\mathbf{P}_{0}\theta+(1-\alpha)\theta^{t}\mathbf{P}_{2}\theta$ $=\theta^{t}(\alpha\mathbf{P}_{0}+(1-\alpha)\mathbf{P}_{2})\theta$